# thanks: Corresponding author

### v2

@@ III.1 Geometry of the black hole-observer system
To facilitate subsequent discussion, it is convenient to partition the BH’s accretion disk along $OX$-axis by the $XOY^{\prime\prime}$ plane (the green plane in Fig. 2) into two distinct semi-disks: the co-side semi-disk (located on the same side as the observer, to the right of the $XOY^{\prime\prime}$ plane) and the counter-side semi-disk (on the opposite side, to the left of the $XOY^{\prime\prime}$ plane).
@@ III.2 Classification of photon trajectories
When photons arrive at the plane $XOY^{\prime\prime}$ from a distant observer, they undergo an azimuthal angle change of $\pi/2$. In Fig. 3, we single out the trajectory plane (the pink plane) in Fig. 2 and take the case where photons intersect each semi-disk once as an example, in order to show the dependence of the total changes $\phi$ on these two semi-disks. The two points $P_{2}$ and $P_{1}$ represent the intersection points of photons with the counter-side and co-side semi-disks, respectively. It is easy to see that from Fig. 3 one only needs to calculate the angle $\lambda$ between the plane $XOY^{\prime\prime}$ and the accretion disk within this trajectory plane to determine the total azimuthal angle of photons reaching the two semi-disks, and thus obtain the total change $\phi=\pi/2\pm\lambda$, where $+$ for the counter-side semi-disk while -- for the co-side semi-disk. Based on the geometric relations in Fig. 2, we get
@@
In Figs. 5 and 6, we present the variations of orbit numbers $n_{1}$ and $n_{2}$ with impact parameter $b$ for different values of $\omega$ and $\eta$, respectively. In Fig. 5, with the polar angle fixed at $\eta=\pi/2$, the impact parameter range corresponding to the lensed emission region of the counter-side semi-disk gradually decreases with increasing viewing inclination, while that of the co-side semi-disk shows significant enlargement. Figure 6 demonstrates similar trends at fixed inclination $\omega=17\pi/36$ across various polar angles $\eta$. Additionally, we find the orbit numbers remain invariant under inclination changes when $\eta=0$.
It is crucial to emphasize that for a given polar angle $\eta$ in the observer’s plane, the counter-side and co-side semi-disks are imaged in the upper and lower half-planes of the $O^{\prime}X^{\prime}$ axis in the observer’s coordinate system, respectively, as shown in Fig. 4. Moreover, the observed image exhibits exact symmetry about the $O^{\prime}Y^{\prime}$ axis. This symmetry reduces our analysis to the range $20\leq\eta\leq\pi/$ 2 while fully characterizing the entire image. All subsequent $\eta$ values are implicitly constrained to this angular domain without repeated specification.
@@ III.3 Transfer functions
Figures 7 and 8 respectively present the transfer functions for the counter-side semi-disk and the first three transfer functions for the co-side semi-disk, with fixed $\eta=\pi/2$ and $\omega=17\pi/36$. Here, black, orange, and red curves correspond to the cases of $m=1$, $m=2$, and $m=3$, respectively. The results align with prior findings on orbit numbers: as inclination $\omega$ (or polar angle $\eta$) increases, the impact parameter ranges for the second ($m=2$) and third ($m=3$) transfer functions of the counter-side semi-disk shrink, while those of the co-side semi-disk expand. Moreover, the slope of the first ($m=1$) transfer function grows progressively with increasing $\omega$ and $\eta$, accompanied by a reduction in its impact parameter range, which indicates a contraction of the direct emission region on the observer’s plane with a bigger $\omega$.
@@ IV Optical appearance of Schwarzschild black hole with a optically thin disk at different inclinations
Combining Eqs. (12), (13), and (16), we calculate the received intensity on the observer’s plane for different $\eta$ at inclination angle $\omega$ with values $\pi/4$ and $3617\pi/$ 36 for the emission intensity characterized by the Model-I, and plotted the results in Fig. 10 and 11. From the figures, we observe that for the counter-side semi-disk, the number of intensity peaks in $I_{\rm obs1}$ shows no significant variation with polar angle $\eta$, while the distance between adjacent peaks displays only minor adjustments. Meanwhile, increasing the inclination angle reduces both the intensity at fixed $\eta$ and the peak separation distance. In contrast, the co-side intensity $I_{\rm obs2}$ exhibits more complex intensity profiles, developing additional emission peaks. For Model-II and Model-III, we briefly present the received intensities for select polar angles at inclination angles $\omega=\pi/4$ and $\omega=17\pi/36$ in Appendix A.
Figure 11: For Model-I at $3617\pi/$ 36 inclination, the counter-side and co-side intensities are shown for varying polar angles $\eta$. The upper and lower rows show $I_{\text{obs1}}$ and $I_{\text{obs2}}$ respectively.
(a) Model-I:$\pi/36$
(b) Model-II:$\pi/36$
(c) Model-III:$\pi/36$
(d) Model-I:$18017\pi/$ 180
(e) Model-II:$18017\pi/$ 180
(f) Model-III:$18017\pi/$ 180
(g) Model-I:$\pi/4$
(h) Model-II:$\pi/4$
(i) Model-III:$\pi/4$
(j) Model-I:$17\pi/36$
(k) Model-II:$17\pi/36$
(l) Model-III:$17\pi/36$
Figure 12: Optical appearances of a Schwarzschild BH illuminated by three types of geometrically and optically thin accretion disks at varying inclination angles. Columns 1-3 display images for Model-I, Model-II, and Model-III at inclinations $\omega=\pi/36$, $18017\pi/$ 180, $\pi/4$, and $3617\pi/$ 36, respectively.
We then compute the received intensity for multiple polar angles at each inclination, and combine these results with the symmetry properties to project the intensity distribution onto a two-dimensional plane. This yields the final optical appearance of a thin-disk-illuminated Schwarzschild BH at various inclinations. As shown in Fig. 12, we present the optical appearances for three thin disk models at inclination angles of $\pi/36$, $18017\pi/$ 180, $\pi/4$, and $3617\pi/$ 36. And we find that as the inclination angle $\omega$ increases, the partial bright rings in the optical appearance of thin accretion disks for all three models become progressively compressed and deviate from circularity.
@@ V Optical appearance of Schwarzschild black hole with a optically thick disk at different inclinations
We now examine the optical appearance of a Schwarzschild BH surrounded by optically thick versions of the three accretion disk models. Unlike the optically thin case, the optically thick disk obstructs photon propagation, preventing some rays from reaching the observer and resulting in darker images. Therefore, to determine the final appearance, the spatial region of the thick accretion disk has to be specified. Based on the emission profiles of the three models, we define the disk’s inner boundaries at locations where the emission intensity drops to zero for the three emission models. Furthermore, due to the rapid emission decay in all models, we consider finite outer boundaries for these accretion disks. In what follows, we consider the regions of the optically thick disks with radius $r$ as
@@
The analysis from previous sections remains applicable here. For Model-I, the impact parameter range in the observer’s plane, corresponding to the first transfer function $r_{1}(b)$ and $R_{1}(b)$ in $[r_{\rm isco},12]$, has no contributions from lensed or photon ring emission due to obstruction by the thick disk. However, outside this impact parameter range, the lensed and photon ring emission will contribute to the image intensity. Similar results hold for Model-II and Model-III. We therefore follow the earlier methodology, using Eq. (16) to calculate the received intensity for each transfer function.
In Fig. 13, we take $\omega=17\pi/36$ as an example to demonstrate the optical appearance of a Schwarzschild BH illuminated by thick accretion disks under the three models, with direct comparison to the thin-disk cases. The top row of Fig. 13 displays the images for the three thin-disk models, while the bottom row shows their thick-disk counterparts. Two key effects are evident: (i) the bright rings disappear in the thick-disk images due to disk obscuration, and (ii) the overall intensities of the thick disks is approximately $20\%$$40\%$ lower than those of the thin disks.
(a) Model-I:$17\pi/36$ (thin)
(b) Model-II:$17\pi/36$ (thin)
(c) Model-III:$17\pi/36$ (thin)
(d) Model-I:$17\pi/36$ (thick)
(e) Model-II:$17\pi/36$ (thick)
(f) Model-III:$17\pi/36$ (thick)
Figure 13: Optical appearances of the Schwarzschild BH illuminated by the three accretion disk models at an inclination angle of $\omega=17\pi/36$. The top row displays images for the three optically thin disk models, while the bottom row shows their optically thick counterparts.
@@ VI Summary
Following the methodology of [31], we categorized light rays intersecting the disk once, twice, and three or more times as direct, lensed, and photon ring trajectories, respectively. We then presented the photon trajectories in Fig. 4 and orbit numbers in Figs. 5 and 6 for both semi-disks at fixed inclination $\omega$ and polar angle $\eta$. The results reveal striking differences between the two semi-disks: the lensed emission region is larger on the co-side, and its size increases markedly with inclination, while the corresponding region on the counter-side shrinks under the same conditions.
@@
To establish the relationship between the accretion disk and points on the observer’s plane, we introduced transfer functions $r_{\rm m}(b)$ in Eq. (12) and $R_{\rm m}(b)$ in Eq. (13) for the counter-side and co-side semi-disks, respectively. These functions characterize the disk intersection positions of light rays received at each point on the observer’s plane. We presented the variation of these transfer functions with inclination angle $\omega$ and polar angle $\eta$ in Figs. 7 and 8. The first ($m=1$) transfer functions for both semi-disks exhibit progressively steeper slopes with increasing $\omega$ and $\eta$, accompanied by a reduction in the corresponding impact parameter ranges. In contrast, the second ($m=2$) and third ($m=3$) transfer functions of both semi-disks exhibit markedly distinct behaviors with increasing $\omega$ and $\eta$. Notably, when both the $\omega$ and $\eta$ approach $\pi/2$, the lensed emission region of the co-side semi-disk dominates the image contribution.
@@
Subsequently, by specifying the emission profiles of the thin disk models and incorporating the transfer functions, we calculated the received intensity at different polar angles $\eta$ for a fixed inclination $\omega$, with partial results presented in Fig. 10 and 11. Our analysis demonstrates that the co-side semi-disk exhibits more complex peak structures in the received intensity across varying polar angles $\eta$, whereas the counter-side semi-disk shows comparatively minor intensity variations. Moreover, we calculated the received intensity for each model at various $\omega$ and projected these results onto a two-dimensional plane to obtain the optical appearances of the three thin disk models under different inclinations, as shown in Fig. 12.
@@
Finally, we extended our discussion to optically thick versions of the three models, comparing their images with the thin disk cases using $\omega=17\pi/36$ as a representative example in Fig. 13. The results demonstrate that the thick disk’s obscuration effect eliminates certain bright rings and reduces the overall intensity by approximately $20\%$$40\%$ compared to thin disks.
In conclusion, this study extends the face-on imaging analysis of three thin accretion disk models to general inclination angles, better matching real astronomical observation conditions. We have obtained optical images for both thin and thick disk scenarios across various inclinations, enriching theoretical understanding and providing valuable references for future observational studies.
@@ Acknowledgements.
This work is supported in part by NSFC Grant No. 12165005.

### v3

@@ III.1 Geometry of the black hole-observer system
Figure 3: Schematic diagram of the co-side and counter-side semi-disks, where $O$ and $O^{\prime}$ denote the positions of the BH and the observer, respectively. The dashed lines $NY^{\prime\prime}$ and $O^{\prime}Y^{\prime}$ represent the intersections of the plane $OO^{\prime}Y^{\prime}$ with the planes $XOY^{\prime\prime}$ and $X^{\prime}O^{\prime}Y^{\prime}$ from Fig. 2, respectively.
To establish a one-to-one correspondence between all points on the accretion disk and those on the observer’s plane, it is convenient to partition the BH’s accretion disk along the $OX$-axis by the $XOY^{\prime\prime}$ plane (the green plane in Fig. 2) into two distinct semi-disks: the co-side semi-disk (located on the same side as the observer, to the right of the $XOY^{\prime\prime}$ plane) and the counter-side semi-disk (on the opposite side, to the left of the $XOY^{\prime\prime}$ plane). Figure 3 shows a schematic diagram of the cross-section formed by intersecting the accretion disk with a plane (a dash line) passing through both the BH and the observer, clearly illustrating the relative positions of the co-side and counter-side semi-disks.
@@ III.2 Classification of photon trajectories
When photons arrive at the plane $XOY^{\prime\prime}$ from a distant observer, they undergo an azimuthal angle change of $\pi/2$. In Fig. 4, we single out the trajectory plane (the pink plane) in Fig. 2 and take the case where photons intersect each semi-disk once as an example, in order to show the dependence of the total changes $\phi$ on these two semi-disks. The two points $P_{2}$ and $P_{1}$ represent the intersection points of photons with the counter-side and co-side semi-disks, respectively. It is easy to see that from Fig. 4 one only needs to calculate the angle $\lambda$ between the plane $XOY^{\prime\prime}$ and the accretion disk within this trajectory plane to determine the total azimuthal angle of photons reaching the two semi-disks, and thus obtain the total change $\phi=\pi/2\pm\lambda$, where ++ for the counter-side semi-disk while - for the co-side semi-disk. Based on the geometric relations in Fig. 2, we get
@@
Through further generalization, we find that for a light ray emitted backward from the observer, which initially intersects the counter-side semi-disk and subsequently experiences multiple intersections with the accretion disk, the azimuthal angle accumulated at each intersection can be expressed as
$\phi=\frac{2m-1}{2}\pi+\arctan{({\tan{\omega}\sin{\eta}})},\;m=1,2,3,\cdots,$ (12)
where $m$ denotes the $m$-th intersection with the accretion disk. Analogously, for the co-side semi-disk, the corresponding azimuthal angle reads
$\phi=\frac{2m-1}{2}\pi-\arctan{({\tan{\omega}\sin{\eta}})},\;m=1,2,3,\cdots.$ (13)
Hence, for a given $m$-th intersection, the the azimuthal angles of the two distinct panels differ. This difference serves as a more significant rationale for partitioning the accretion disk into two panels.
@@
In Figs. 6 and 7, using Eqs. (10), (12) and (13), we present the variations of orbit numbers $n_{1}$ and $n_{2}$ with impact parameter $b$ for different values of $\omega$ and $\eta$, respectively. In Fig. 6, with the polar angle fixed at $\eta=\pi/2$, the impact parameter range corresponding to the lensed emission region of the counter-side semi-disk gradually decreases with increasing viewing inclination, while that of the co-side semi-disk shows significant enlargement. Figure 7 demonstrates similar trends at fixed inclination $\omega=17\pi/36$ across various polar angles $\eta$. Additionally, we find the orbit numbers remain invariant under inclination changes when $\eta=0$.
It is crucial to emphasize that for a given polar angle $\eta$ in the observer’s plane, the counter-side and co-side semi-disks are imaged in the upper and lower half-planes of the $O^{\prime}X^{\prime}$ axis in the observer’s coordinate system, respectively, as shown in Fig. 5. Moreover, the observed image exhibits exact symmetry about the $O^{\prime}Y^{\prime}$ axis. This symmetry reduces our analysis to the range $0\leq\eta\leq\pi/2$ while fully characterizing the entire image. All subsequent $\eta$ values are implicitly constrained to this angular domain without repeated specification.
@@ III.3 Transfer functions
Figures 8 and 9 respectively present the transfer functions for the counter-side semi-disk and the first three transfer functions for the co-side semi-disk, with fixed $\eta=\pi/2$ and $\omega=17\pi/36$. Here, black, orange, and red curves correspond to the cases of $m=1$, $m=2$, and $m=3$, respectively. The results align with prior findings on orbit numbers: as inclination $\omega$ (or polar angle $\eta$) increases, the impact parameter ranges for the second ($m=2$) and third ($m=3$) transfer functions of the counter-side semi-disk shrink, while those of the co-side semi-disk expand. Moreover, the slope of the first ($m=1$) transfer function grows progressively with increasing $\omega$ and $\eta$, accompanied by a reduction in its impact parameter range, which indicates a contraction of the direct emission region on the observer’s plane with a bigger $\omega$.
@@ IV Optical appearance of Schwarzschild black hole with a optically thin disk at different inclinations
$\text{Model-I:}\quad I_{\rm em}(r):=\begin{cases}I_{0}\left[\frac{1}{r-(r_{\rm isco}-1)}\right]^{2},&\hskip 17.07182ptr>r_{\rm isco}\\ 0,&\hskip 17.07182ptr\leqslant r_{\rm isco}\end{cases},$ (19)
$\text{Model-II:}\quad I_{\rm em}(r):=\begin{cases}I_{0}\left[\frac{1}{r-(r_{\rm ph}-1)}\right]^{3},&\hskip 25.6073ptr>r_{\rm ph}\\ 0,&\hskip 25.6073ptr\leqslant r_{\rm ph}\end{cases},$ (20)
$\text{Model-III:}\quad I_{\rm em}(r):=\begin{cases}I_{0}\frac{\frac{\pi}{2}-\arctan[r-(r_{\rm isco}-1)]}{\frac{\pi}{2}-\arctan[r_{\rm h}-(r_{\rm isco}-1)]},&\hskip 2.84544ptr>r_{\rm h}\\ 0,&\hskip 2.84544ptr\leqslant r_{\rm h}\end{cases},$ (21)
Combining Eqs. (14), (15), and (18), we calculate the received intensity on the observer’s plane for different $\eta$ at inclination angle $\omega$ with values $\pi/4$ and $17\pi/36$ for the emission intensity characterized by the Model-I, and plotted the results in Fig. 11 and 12. From the figures, we observe that for the counter-side semi-disk, the number of intensity peaks in $I_{\rm obs1}$ shows no significant variation with polar angle $\eta$, while the distance between adjacent peaks displays only minor adjustments. Meanwhile, increasing the inclination angle reduces both the intensity at fixed $\eta$ and the peak separation distance. In contrast, the co-side intensity $I_{\rm obs2}$ exhibits more complex intensity profiles, developing additional emission peaks. For Model-II and Model-III, we briefly present the received intensities for select polar angles at inclination angles $\omega=\pi/4$ and $\omega=17\pi/36$ in Appendix A.
Figure 12: For Model-I at $17\pi/36$ inclination, the counter-side and co-side intensities are shown for varying polar angles $\eta$. The upper and lower rows show $I_{\text{obs1}}$ and $I_{\text{obs2}}$ respectively.
@@
[…]
Figure 13: Optical appearances of a Schwarzschild BH illuminated by three types of geometrically and optically thin accretion disks at varying inclination angles. Columns 1-3 display images for Model-I, Model-II, and Model-III at inclinations $\omega=\pi/36$, $17\pi/180$, $\pi/4$, and $17\pi/36$, respectively.
We then compute the received intensity for multiple polar angles at each inclination, and combine these results with the symmetry properties to project the intensity distribution onto a two-dimensional plane. This yields the final optical appearance of a thin-disk-illuminated Schwarzschild BH at various inclinations. As shown in Fig. 13, we present the optical appearances for three thin disk models at inclination angles of $\pi/36$, $17\pi/180$, $\pi/4$, and $17\pi/36$. And we find that as the inclination angle $\omega$ increases, the partial bright rings in the optical appearance of thin accretion disks for all three models become progressively compressed and deviate from circularity.
@@ V Optical appearance of Schwarzschild black hole with a optically thick disk at different inclinations
For accretion disks, when scattering of photons by the disk is neglected, the absorption degree of radiation is typically characterized by the “true” absorption optical depth parameter $\bar{\tau}$, defined as [19, 61]
$\bar{\tau}=\sigma_{\rm ff}U_{0}.$ (22)
Here $\sigma_{\rm ff}$ and $U_{0}$ denote the free-free absorption coefficient and the surface density of the disk, respectively. In general, an optical depth $\bar{\tau}>1$ signifies strong absorption by the disk, indicating that the disk is opaque (i.e., optically thick). Guided by this definition, we next investigate the optical appearance of a Schwarzschild BH surrounded by optically thick versions of the three accretion disk models. Unlike the optically thin case, the optically thick disk obstructs photon propagation, preventing some rays from reaching the observer and resulting in darker images [19]. Therefore, to determine the final appearance, the spatial region of the thick accretion disk has to be specified. Based on the emission profiles of the three models, we define the disk’s inner boundaries at locations where the emission intensity drops to zero for the three emission models. Furthermore, due to the rapid emission decay in all models, we consider finite outer boundaries for these accretion disks. In what follows, we consider the regions of the optically thick disks with radius $r$ as
@@
The analysis from previous sections remains applicable here. For Model-I, the impact parameter range in the observer’s plane, corresponding to the first transfer function $r_{1}(b)$ and $R_{1}(b)$ in $r_{\rm isco},12]$, has no contributions from lensed or photon ring emission due to obstruction by the thick disk. However, outside this impact parameter range, the lensed and photon ring emission will contribute to the image intensity. Similar results hold for Model-II and Model-III. We therefore follow the earlier methodology, using Eq. (18) to calculate the received intensity for each transfer function.
In Fig. 14, we take $\omega=17\pi/36$ as an example to demonstrate the optical appearance of a Schwarzschild BH illuminated by thick accretion disks under the three models, with direct comparison to the thin-disk cases. The top row of Fig. 14 displays the images for the three thin-disk models, while the bottom row shows their thick-disk counterparts. Two key effects are evident: (i) the bright rings disappear in the thick-disk images due to disk obscuration, and (ii) the overall intensities of the thick disks are lower than those of the thin disks.
[…]
Figure 14: Optical appearances of a Schwarzschild BH illuminated by the three accretion disk models at an inclination angle of $\omega=17\pi/36$. The top row displays images for the three optically thin disk models, while the bottom row shows their optically thick counterparts.
@@ VI Summary
Following the methodology of [31], we categorized light rays intersecting the disk once, twice, and three or more times as direct, lensed, and photon ring trajectories, respectively. We then presented the photon trajectories in Fig. 5 and orbit numbers in Figs. 6 and 7 for both semi-disks at fixed inclination $\omega$ and polar angle $\eta$. The results reveal striking differences between the two semi-disks: the lensed emission region is larger on the co-side, and its size increases markedly with inclination, while the corresponding region on the counter-side shrinks under the same conditions.
@@
To establish the relationship between the accretion disk and points on the observer’s plane, we introduced transfer functions $r_{\rm m}(b)$ in Eq. (14) and $R_{\rm m}(b)$ in Eq. (15) for the counter-side and co-side semi-disks, respectively. These functions characterize the disk intersection positions of light rays received at each point on the observer’s plane. We presented the variation of these transfer functions with inclination angle $\omega$ and polar angle $\eta$ in Figs. 8 and 9. The first ($m=1$) transfer functions for both semi-disks exhibit progressively steeper slopes with increasing $\omega$ and $\eta$, accompanied by a reduction in the corresponding impact parameter ranges. In contrast, the second ($m=2$) and third ($m=3$) transfer functions of both semi-disks exhibit markedly distinct behaviors with increasing $\omega$ and $\eta$. Notably, when both the $\omega$ and $\eta$ approach $\pi/2$, the lensed emission region of the co-side semi-disk dominates the image contribution.
@@
Subsequently, by specifying the emission profiles of the thin disk models and incorporating the transfer functions, we calculated the received intensity at different polar angles $\eta$ for a fixed inclination $\omega$, with partial results presented in Fig. 11 and 12. Our analysis demonstrates that the co-side semi-disk exhibits more complex peak structures in the received intensity across varying polar angles $\eta$, whereas the counter-side semi-disk shows comparatively minor intensity variations. Moreover, we calculated the received intensity for each model at various $\omega$ and projected these results onto a two-dimensional plane to obtain the optical appearances of the three thin disk models under different inclinations, as shown in Fig. 13.
@@
Finally, we extended our discussion to optically thick versions of the three models, comparing their images with the thin disk cases using $\omega=17\pi/36$ as a representative example in Fig. 14. The results demonstrate that, compared to thin disks, the thick disk’s obscuration effect eliminates certain bright rings and reduces the overall intensity.
In conclusion, this study extends the face-on imaging analysis of three thin accretion disk models to general inclination angles, better matching real astronomical observation conditions. We have obtained optical images for both thin and thick disk scenarios across various inclinations, enriching theoretical understanding and providing valuable references for future observational studies. The framework developed in this paper can be applied to a broader range of accretion disk models, though with increased complexity. For example, extending from geometrically thin to thick disks requires replacing the two-dimensional disk surface with a three-dimensional emitting structure. This not only introduces additional radiation from the sides of the disk but also enables the disk to obstruct the propagation of light, thereby altering the visibility of characteristic structures such as photon rings. Furthermore, while the current method is well-suited for static, spherical BHs, extending it to the Kerr spacetime (which describes rotating BHs) remains a challenge due to the additional complexity introduced by rotation. The axisymmetric and complex structure of such spacetimes breaks existing simplifications and increases the computational difficulty of ray tracing and image synthesis. The primary objective of our future work is to address this challenge by extending the present framework, with the goal of producing more comprehensive and observationally faithful BH optical images.
@@ Acknowledgements.
This work is supported in part by NSFC Grants No. 12165005 and No. 11961131013.